# SlipStream: Adapting Pipelines for Distributed Training of Large DNNs Amid Failures

### v1

@@ 4.3. Implementation
We implemented SlipStream on top of DeepSpeed (Rasley et al., 2020) and made the following additions to DeepSpeed to support SlipStream.
@@
Rerouting micro-batches to data parallel peers. To handle the dynamic rerouting of micro-batches following node failures, we introduce two new and complementary communication operators: ReRouteAct and ReRouteGrad. Positioned at the end of a pipeline stage, ReRouteAct normally acts as a pass-through, transmitting computed intermediate activations to the next stage within the same data parallel pipeline. Conversely, ReRouteGrad, located at the beginning of a stage, typically forwards gradients backward through the pipeline. When a subsequent stage fails, ReRouteAct redistributes the micro-batches across the remaining peers in a round-robin fashion, while ReRouteGrad adjusts the gradient distribution to ensure that both the forward and backward processes of a micro-batch are handled by the same peer. This strategy maintains operational continuity in SlipStream under fault conditions, mirroring fault-free execution semantics. These operators are integrated as pipeline instructions within the DeepSpeed execution engine.
@@
Decoupling Back Propagation in DeepSpeed. Our implementation of decoupled backpropagation within DeepSpeed centers around intercepting the weight gradient computations traditionally performed during the backpropagation phase. These computations are temporarily held in a newly created in-memory structure known as WeightGradStore, enabling SlipStream to defer weight gradient computation. To facilitate this, two new pipeline instructions have been introduced into DeepSpeed’s execution engine: InputBackwardPass and WeightBackwardPass. These instructions are incorporated into SlipStream’s execution plan and are then processed by DeepSpeed’s execution scheduler, effectively managing the distinct phases of backpropagation.
@@
Bypassing Optimizer Synchronizations. In DeepSpeed, each training iteration concludes with an all-reduce collective to synchronize gradients across data parallel peers, followed by a series of validation checks at each pipeline stage to ensure numerical stability, such as monitoring for gradient overflow. If any stage identifies an issue that could compromise numerical stability, the optimizer step for that iteration is omitted to avoid exacerbating the problem. Upon successful validation, the optimizer step is executed for each stage, and the training progresses to the next iteration in a synchronized manner across all stages.
@@
Table 1. Normalized training throughput with increasing failure rates. Fault-free training represents $1.00$, hence higher is better. Bamboo ran out of memory for GPT-3 3.35B and 6.7B.
@@
Bamboo (Thorpe et al., 2023) 0.71 0.69 0.55 OOM OOM OOM OOM OOM OOM
Oobleck (Jang et al., 2023) 0.99 0.92 0.71 0.98 0.90 0.66 0.93 0.87 0.52
SlipStream 0.99 0.92 0.81 0.98 0.95 0.85 0.97 0.91 0.66
In contrast, SlipStream employs a staggered timing approach for the optimizer steps across different pipeline stages to enhance scheduling efficiency. This method of staggering is not conducive to cross-stage synchronization for validating numerical stability prior to executing the optimizer. To accommodate staggered execution while maintaining validation integrity, numerical validations are shifted from a pre-step to a post-step process. Following the all-reduce collective, each stage immediately conducts its own local validation checks but does not wait for downstream stages to complete their validations. Each stage proceeds with its optimizer step based on its individual validation results and those of preceding stages. If a downstream stage fails its validation, a rollback is initiated across all stages before advancing to the next iteration. Notably, for many optimizers, including the widely-used AdamW (Loshchilov and Hutter, 2017), such rollbacks do not entail additional memory costs because the operations involved are arithmetically reversible. This adjustment allows SlipStream to retain the advantages of staggered operations while protecting the training process from potential instabilities.
@@ 5. Evaluation
Figure 8. Training throughput, normalized to the fault-free case, for the GPT-3 Medium and GPT-3 6.7B models over the GCP trace. In 8b and 8c, the dashed lines represent the average training throughput achieved by each system within the 6h period.
@@ 5.1. Experimental Setup
Cluster Setup. We performed real-world experiments on a cluster of 32 NVIDIA A100 80GB GPUs in Azure. Specifically, we used four Standard_NC96ads_A100_v4 instances, each with 8 GPUs, 96 vCPUs, and 880 GB of host memory. Each node has an 600 GB/s NVLink intra-node interconnect, as well as a 640 Gbps inter-node interconnect over 8 NICs. We performed scalability experiments using a simulator, which we present and validate in Section 5.3.
Baselines. We compared SlipStream to the two state-of-the-art baselines discussed in Section 2.3 – Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023) – with all of their respective optimizations enabled. We report the fault-free throughput achieved by DeepSpeed (Rasley et al., 2020) using an 1F1B pipeline schedule (Narayanan et al., 2019). All baselines train on the same Azure cluster.
Workloads. We evaluated all systems using the Megatron (Shoeybi et al., 2019) implementation of GPT-3 (Brown et al., 2020). We used three GPT-3 sizes: Medium (350M), 3.35B, and 6.7B. We used (PP, DP) degrees of (2, 16), (4, 8), and (8, 4) for the Medium, 3.35B, and 6.7B models, respectively. All models used a TP degree of 1. We trained on wikitext (Merity et al., 2016) using (batch, micro-batch) size of (8192, 8), (1024, 1), and (1024, 1) for the Medium, 3.35B, and 6.7B models, respectively. Unless otherwise noted, all real-world experiments ran for 6 hours.
@@ 5.2. Training Throughput Under Failures
How well does SlipStream handle failures compared to baselines?
For each 32-GPU training job on Azure, we instrumented a random GPU to fail at regular and increasing intervals: 6 hours, 2 hours, and 30 minutes, which is representative of real-world failure rates (Jeon et al., 2019; Weng et al., 2022), starting at the beginning of the training job ($t=0$). Note that we treat each individual GPU as a worker. Table 1 shows the average training throughput achieved by each system across models and failure rates normalized to the fault-free throughput achieved by DeepSpeed ($1.00$). Note that GPUs are not replaced after failure, hence the total number of workers monotonically declines throughout each experiment. With a failure every 30 minutes, the percentage of functional GPUs by the end of the 6 hour run is 62.5%. Hence, we expect the average throughput with a failure every 30min to be relatively modest compared to the fault-free case.
Because Bamboo relies on redundant computation, it rapidly exhausts GPU memory with its additional copy of model states and activations. Hence, Bamboo fails to train the two larger models. Even for the smallest model, redundant computation introduces significant throughput degradation. A single GPU failure (3% of total GPUs) in the 6h job causes a 29% throughput drop. Bamboo actually penalizes fault-free execution by nearly the same percentage compared to DeepSpeed. Oobleck can handle all model sizes and achieves significant throughput improvements over Bamboo. However, as the frequency of failures and model size increase, Oobleck’s performance degrades. As failures increase, Oobleck’s heterogeneous pipelines become more imbalanced, degrading the throughput of the synchronous training job. As the model size increases, the latency of reconfiguring all of the stages in the pipeline with the failed GPU gets higher.
SlipStream is able to efficiently continue training under failures, out-performing or matching the throughput of both baselines across all models and failure rates. Adaptive Pipelining, Decoupled BackProp, and Staggered Optimizer are effective in optimizing the use of compute and memory resources across peers of failed GPUs. Moreover, SlipStream recovers from a failure quickly as it migrates model parameters to only a single GPU during normalization (see Section 4.2.1). Under frequent failures (every 30m) and no repairs, SlipStream achieves 85% and 66% of the fault-free training throughput for GPT-3 3.35B and 6.7B respectively. SlipStream out-performs Oobleck and Bamboo by up to 1.29×1.29$\times$1.29 × and 1.46×1.46$\times$1.46 ×, respectively for these experiments. Note that, unlike Bamboo, both SlipStream and Oobleck introduce no penalty over DeepSpeed for fault-free execution.
What is SlipStream’s throughput advantage in an realistic training scenario?
Next, we evaluated a realistic training scenario where GPUs fail and re-join based on a real-world trace. We used a trace of node failures/repairs from GCP instances used by Bamboo (Thorpe et al., 2023). We observed similar results for Bamboo’s AWS trace, but omitted the plots due to space constraints. We replayed the events in the GCP trace over a 6h training run on our Azure cluster. At the time of the experiment, we had 24 GPUs available (instead of 32 GPUs). Figure 8 shows the number of available GPUs over time according to the trace (max 24, min 15).
@@
Figure 8 also plots the training throughput achieved by Bamboo, Oobleck, and SlipStream replaying the trace for a GPT-3 Medium and GPT-3 6.7B training job. Solid lines represent instantaneous training throughput, while dashed lines show the average training through for each system over the 6h period. Bamboo cannot train GPT-3 6.7B due to memory pressure. SlipStream achieves a 1.64×1.64$\times$1.64 × higher performance (average throughput) compared to Bamboo on GPT-3 Medium and a 1.46×1.46$\times$1.46 × higher performance than Oobleck on GPT-3 6.7B. As discussed above, Bamboo suffers from the overheads of redundant computations. Oobleck requires large stalls due to parameter re-shuffling both when GPUs fail and when they rejoin, causing frequent and large drops in throughput. Meanwhile, during periods of stability, Oobleck suffers from imbalanced heterogeneous pipelines as discussed above. SlipStream provides the highest and most consistent training throughput across all portions of the trace.
5.3. SlipStream Scalability
@@
Since we do not have access to a cluster of thousands of GPUs, we built a simulator that can calculate the training throughput of a model and cluster configuration, given an execution plan. The simulator uses real-world profiled statistics for each pipeline operation of the respective model. To demonstrate that the simulator yields accurate results, Table 2 shows the gap between the simulated throughput and measured real-world througpught on Azure across different failure rates for the three GPT-3 models. The maximum difference is $5.98\%$. These differences arise due to minor fluctuations in the execution time of NCCL collectives, which do not cause impactful discrepancies when evaluating SlipStream.
How well does SlipStream scale to large clusters and models?
Using the simulator, we explored how well SlipStream handles failures for large models training on larger clusters. Figure 9 shows of the simulated throughput achieved by SlipStream across four different GPT-3 models with up to 145.6B parameters, normalized to the simulated fault-free throughput achieved by DeepSpeed’s 1F1B schedule. We also show the fault-scaled throughput, which is obtained by multiplying the fault-free throughput by the percent of non-failed GPUs. Note that we report the steady-state throughput with 1%, 5%, and 10% failed GPUs, not the throughput as some number of GPUs fails and rejoins. We simulated a cluster of (# GPUs, # stages per pipeline, and # pipelines) = (256, 8, 32), (512, 16, 32), (1024, 32, 32) and (1536, 64, 24) for the 18.4B, 39.1B, 76.1B, and 145.6B model, respectively, which reflect the larger cluster sizes needed to train large models.
@@
Figure 9. Simulated throughput of SlipStream as model size increases, normalized to the simulated fault-free 1F1B throughput. The Fault-Scaled throughput is the fault-free throughput scaled by the percent of non-failed GPUs.
@@
SlipStream is effective at achieving high throughput across failure rates and model sizes. For the case of 1% failures (higher reliability), SlipStream handles failures at or better than the fault-scaled throughput across all models and cluster sizes, often with zero degradation. The bubbles in peer workers are more than sufficient to efficiently execute the work for failed GPUs using SlipStream’s optimizations (see Figure 6). As the number of failed GPUs increases, SlipStream is still able to sustain high throughput. At $5\%$ GPU failures, SlipStream achieves comparable performance to the fault-scaled throughput. Even at a large $10\%$ failure rate (e.g., 154 failed GPUs for GPT-3 145.6B), SlipStream enables clusters to continue training while only observing between $0.5\%$ and $11.5\%$ degradation to the fault-scaled case. This suggests that SlipStream will be effective for training jobs in supercomputing-scale clusters with dynamic resource availability (Jiang et al., 2024; He et al., 2023; Zu et al., 2024).
5.4. SlipStream Performance Breakdown
@@
Figure 10. The contribution of the three optimization techniques to SlipStream’s training throughput.
@@
Figure 11. Memory utilization for SlipStream across pipeline stages for GPT-3 6.7B with 30m failure rates compared to DeepSpeed (failure-free). The dashed line shows the 80GB memory capacity of the A100 GPU.
How does each of SlipStream’s techniques contribute to performance?
@@
To understand the benefit (and necessity) of each of the SlipStream techniques described in Section 3, we performed an ablation study by repeating the real-world experiment on the Azure cluster with the GPT-3 models and 30min failure frequency. We modified the MILP formulation in the Planner to progressively enable optimizations. Figure 10 shows the throughput achieved in each step of the study, normalized to the fault-free throughput.
@@
Adaptive Pipelining allows SlipStream to continue training in the face of failures, but suffers from a large throughput degradation due to overheads caused by extra work in the peers of failed workers. By adding Decoupled BackProp, SlipStream boosts the training throughput by 63% to 118% by using bubbles in the cool-down phase of the pipeline schedule to mask this overhead (see Figure 3(b)). The addition of the Staggered Optimizer further improves throughput by 7% to 11% up to $11\%$ as it allows the use of bubbles from the warm-up phase of the pipeline schedule.
Can SlipStream efficiently exploit unused GPU memory?
@@
SlipStream makes the key insight that GPU memory capacity is not uniformly utilized throughout the pipelined execution. Later pipeline stages inherently use less GPU memory. Decoupled BackProp utilizes this free GPU memory as a buffer to defer $B_{weight}$ computations (see Figure 5). To explore this in more detail, we analyzed the real-world training of the GPT-3 6.7B model with a 30m failure interval as in Section 5.2. Figure 11 shows the peak GPU memory utilization across the 8 stages of a pipelined execution, compared to the memory utilization of the failure-free DeepSpeed baseline.
@@
DeepSpeed, and any system that uses a pipeline schedule similar to 1F1B, does not fully utilize GPU memory, especially in later pipeline stages (Narayanan et al., 2019). This is because GPU workers that execute earlier pipeline stages must store a larger number of intermediate results, as shown in Figure 3. Figure 11 shows how SlipStream exploits this opportunity and uniformly utilizes nearly all GPU memory to optimize pipelined execution. The MILP formulation is able to carefully model and manage the additional memory requirements within GPU memory constraints and avoid memory exhaustion.
@@
Table 3. SlipStream Planner latency (in seconds) to find optimized schedules for up to $25\%$ of failed GPUs.
DP PP Degree
Degree 2 4 8 16 32 64
2 5.24 5.41 15.48 42.73 129.57 459.57
4 7.65 8.62 24.48 59.75 190.54 719.4
8 8.56 9.57 30.74 81.42 294.21 1038.2
16 11.84 13.8 51.53 116.8 452.98 1721.45
32 11.71 14.87 67.24 278.64 942.42 3153.07
What is the overhead of SlipStream’s Planner?
@@
To understand how fast the Planner can generate optimized schedules for large clusters, Table 3 shows the time required to generate all necessary schedules for up to $25\%$ of node failures across various parallelism strategies. Recall that the user can configure a smaller or larger fault-tolerance threshold. The reported latency includes both Planner phases. We report the results using Gurobi MILP solver (Gurobi, 2024) running on an 96-core CPU. We can generate all solutions to support up to $512$ failures in the 2048-GPU case ($DP=32$, $PP=64$) within 3153 seconds (52.5 minutes). Given that the training time for large models on thousands of GPUs is measured in weeks (Smith et al., 2022; Jiang et al., 2024; Naumov et al., 2020), the Planner’s latency is insigificant.

### v2

@@ 5. Implementation
We implemented ReCycle on top of DeepSpeed (Rasley et al., 2020) and made the following additions to DeepSpeed to support ReCycle.
Detecting failures. ReCycle uses existing hardware error detection mechanisms in GPUs (e.g.; SMBPBI APIs in NVIDIA GPUs) and software error detection in runtime systems (e.g. stderr logs in PyTorch). Workers send periodic heartbeats to a central driver, including worker information and hardware statistics from GPUs such as page retirement count, row-remapping stats, ECC correction stats, XID error logs, and stdout/stderr logs (Jiang et al., 2024). If the central driver detects abnormalities, it marks the worker as failed.
@@
Rerouting micro-batches to data parallel peers. To handle the dynamic rerouting of micro-batches following node failures, we introduce two new and complementary communication operators: ReRouteAct and ReRouteGrad. Positioned at the end of a pipeline stage, ReRouteAct normally acts as a pass-through, transmitting computed intermediate activations to the next stage within the same data parallel pipeline. Conversely, ReRouteGrad, located at the beginning of a stage, typically forwards gradients backward through the pipeline. When a subsequent stage fails, ReRouteAct redistributes the micro-batches across the remaining peers in a round-robin fashion, while ReRouteGrad adjusts the gradient distribution to ensure that both the forward and backward processes of a micro-batch are handled by the same peer. This strategy maintains operational continuity in ReCycle under fault conditions, mirroring fault-free execution semantics. These operators are integrated as pipeline instructions within the DeepSpeed execution engine.
@@
Decoupling Back Propagation in DeepSpeed. Our implementation of Decoupled BackProp within DeepSpeed centers around intercepting the weight gradient computations traditionally performed during the backpropagation phase. These computations are temporarily held in a newly created in-memory structure known as WeightGradStore, enabling ReCycle to defer weight gradient computation. To facilitate this, two new pipeline instructions have been introduced into DeepSpeed’s execution engine: InputBackwardPass and WeightBackwardPass. These instructions are incorporated into ReCycle’s execution plan and are then processed by DeepSpeed’s execution scheduler, effectively managing the distinct phases of backpropagation.
@@
Bypassing Optimizer Synchronizations. In DeepSpeed, each training iteration ends with an all-reduce collective to synchronize gradients across data-parallel peers, followed by validation checks at each pipeline stage to ensure numerical stability. If any stage detects a potential issue, the optimizer step for that iteration is skipped to prevent further problems. Upon successful validation, the optimizer step is executed, allowing synchronized progression to the next iteration.
@@
Table 1. Training throughput (samples/sec) with increasing failure frequency, higher is better. Bamboo ran out of memory for GPT-3 3.35B and 6.7B.
@@
Fault-Free DeepSpeed (Rasley et al., 2020) 27.58 14.87 5.33
Bamboo (Thorpe et al., 2023) 19.47 18.98 15.24 OOM OOM OOM OOM OOM OOM
Oobleck (Jang et al., 2023) 27.26 25.37 19.47 14.55 13.44 9.78 4.98 4.65 2.78
ReCycle 27.27 25.42 22.27 14.59 14.17 12.63 5.17 4.85 3.53
In contrast, ReCycle uses a staggered timing approach for optimizer steps across different pipeline stages to improve scheduling efficiency. This method of staggering is not conducive to cross-stage synchronization for validating numerical stability prior to executing the optimizer. To address this, ReCycle shifts numerical validations from a pre-step to a post-step process. After the all-reduce collective, each stage performs its own local validation checks without waiting for downstream stages. Each stage executes its optimizer step based on its own validation results and those of preceding stages. If a downstream stage fails validation, a rollback occurs across all stages before moving to the next iteration. Notably, for many optimizers, including the commonly used AdamW (Loshchilov and Hutter, 2017), these rollbacks incur no additional memory costs due to the arithmetic reversibility of the operations. This adjustment enables ReCycle to leverage staggered operations while safeguarding the training process from potential instabilities.
@@ 6. Evaluation
Figure 9. Training throughput (samples/sec), higher is better, for the GPT-3 Medium and GPT-3 6.7B models over the GCP trace. In 9b and 9c, the dashed lines represent the average training throughput achieved by each system within the 6h period.
@@ 6.1. Experimental Setup
Cluster Setup. We conducted real-world experiments on a 32-GPU cluster featuring NVIDIA A100 80GB GPUs in Azure, utilizing Standard_NC96ads_A100_v4 instances (8 GPUs, 96 vCPUs, and 880 GB memory each). Each node includes a 600 GB/s NVLink intra-node interconnect and a 640 Gbps inter-node interconnect across 8 NICs. Scalability experiments were conducted using a simulator, discussed in Section 6.3.
Baselines. We evaluated ReCycle against two state-of-the-art baselines discussed in Section 2.2.3 – Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023). Both baselines were tested with all their optimizations enabled. Additionally, we report the fault-free throughput achieved by DeepSpeed (Rasley et al., 2020) using a 1F1B pipeline schedule (Narayanan et al., 2019). All experiments were conducted on the same Azure cluster for consistency.
Workloads. We evaluated all systems using the Megatron (Shoeybi et al., 2019) implementation of GPT-3 (Brown et al., 2020), with three model sizes: Medium (350M), 3.35B, and 6.7B. The applied (PP, DP) degrees were (2, 16), (4, 8), and (8, 4), respectively, with a TP degree of 1 for all models. Training was conducted on wikitext (Merity et al., 2016) with batch and micro-batch sizes of (8192, 8) for Medium, and (1024, 1) for both 3.35B and 6.7B models. All real-world experiments, unless stated otherwise, ran for 6 hours with 32 workers.
@@ 6.2. Training Throughput Under Failures
How well does ReCycle handle failures compared to baselines?
We first evaluate the average training throughput of Bamboo, Oobleck and ReCycle on various failure scenarios. We set the frequency of failures from once every 6 hours to once every 30 minutes to cover a wide spectrum of environments (Jeon et al., 2019; Weng et al., 2022). We monotonically reduce the number of available workers without recovery, so the total number of workers steadily decreases over the course of each experiment – for example in the 30m case, only $62.5\%$ of the workers (20 out of 32 workers) remain at the end of training.
Table 1 presents the average throughput for various failure frequencies and model sizes. Bamboo suffers from static overhead due to redundant computations and additional model state copies, which quickly depletes GPU memory. This leads to its inability to train larger models and a significant drop in throughput for even the smallest model—resulting in a 29% reduction in throughput in the 6-hour failure frequency scenario. In contrast, Oobleck effectively manages all model sizes and demonstrates throughput improvements over Bamboo. However, Oobleck’s performance declines with increasing failure frequency and model size due to imbalanced heterogeneous pipelines and higher reconfiguration latency.
ReCycle effectively continues training despite failures, consistently matching or exceeding the throughput of both baselines across all models and failure rates. Unlike Bamboo, the techniques utilized in ReCycle—Adaptive Pipelining, Decoupled BackProp, and Staggered Optimizer—introduce no static overhead, enabling uninterrupted training for larger models and increasing throughput by up to 1.46×1.46$\times$1.46 × for GPT-3 Medium. Additionally, ReCycle has reduced reconfiguration overhead; during failure normalization, it requires parameter migration for at most one GPU, compared to the need for reconfiguring an entire data pipeline with Oobleck. This efficiency results in up to 1.29×1.29$\times$1.29 × higher training throughput.
What is ReCycle’s throughput advantage in dynamic training scenarios?
We next assessed ReCycle’s performance in a dynamic training scenario characterized by GPU failures and re-joins, using a real-world failure trace. This trace, derived from GCP instances utilized by Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023), was replayed over a 6-hour training run on our Azure cluster. While we observed similar outcomes with Bamboo’s AWS trace, we opted to omit those plots due to space limitations. During the experiment, we had 24 GPUs available instead of the originally planned 32. Figure 9a illustrates the fluctuating number of available GPUs over time, ranging from a maximum of 24 to a minimum of 15. Unlike the previous experiment, this scenario frequently saw GPUs being removed and reintroduced into the cluster.
@@
In Figure 9b and 9c, we present the training throughput achieved by Bamboo, Oobleck, and ReCycle while replaying the trace for both GPT-3 Medium and GPT-3 6.7B training jobs. The solid lines indicate instantaneous training throughput, while the dashed lines represent the average training throughput for each system throughout the 6-hour period. Notably, Bamboo is unable to train GPT-3 6.7B due to memory constraints. ReCycle demonstrates a performance increase of 1.64×1.64$\times$1.64 × in average throughput compared to Bamboo on GPT-3 Medium, and a 1.46×1.46$\times$1.46 × improvement over Oobleck on GPT-3 6.7B. As mentioned earlier, Bamboo incurs overhead from redundant computations, while Oobleck experiences significant stalls due to parameter re-shuffling during GPU failures and re-joins, leading to substantial drops in throughput. Additionally, during stable periods, Oobleck grapples with imbalanced heterogeneous pipelines. In contrast, ReCycle consistently delivers the highest and most stable training throughput throughout the entire trace.
6.3. ReCycle Scalability
@@
Due to the unavailability of a cluster with thousands of GPUs, we developed a simulator capable of calculating the training throughput based on a model and cluster configuration, given a specific execution plan. The simulator leverages real-world profiled statistics for each pipeline operation associated with the respective model. To validate the simulator’s accuracy, Table 2 presents the differences between the simulated throughput and the measured real-world throughput on Azure across various failure rates for three GPT-3 models, with a maximum discrepancy of $5.98\%$. These variations are mainly due to minor fluctuations in the execution time of NCCL collectives, but they have little impact on ReCycle’s performance.
How effectively does ReCycle scale to large clusters and models?
Using the simulator, we report the training throughput for GPT models with sizes of 18.4B, 39.1B, 76.1B, and 145.6B across different clusters: (256 GPUs, 8 stages per pipeline, 32 pipelines), (512 GPUs, 16 stages per pipeline, 32 pipelines), (1024 GPUs, 32 stages per pipeline, 32 pipelines), and (1536 GPUs, 64 stages per pipeline, 24 pipelines). Figure 10 presents the throughput normalized to the fault-free throughput achieved by DeepSpeed’s 1F1B schedule. Additionally, we introduce the fault-scaled throughput, calculated by multiplying the fault-free throughput by the percentage of operational GPUs. Notably, we focus on steady-state throughput for scenarios with 1%, 5%, and 10% GPU failures, rather than the throughput during periods of GPU failures and re-joins.
@@
Figure 10. Simulated throughput of ReCycle as model size increases, normalized to the simulated fault-free 1F1B throughput. The Fault-Scaled throughput is the fault-free throughput scaled by the percent of non-failed GPUs.
@@
ReCycle demonstrates a strong capability to maintain high throughput across varying failure rates and model sizes. For a failure rate of 1%, ReCycle manages failures at or better than the fault-scaled throughput for all models and cluster sizes, frequently exhibiting no degradation in performance. The bubbles in peer workers are more than adequate to efficiently handle the workload of failed GPUs utilizing ReCycle’s optimizations (see Figure 6). As the number of failed GPUs increases, ReCycle continues to sustain high throughput levels. At a 5% failure rate, ReCycle’s performance is comparable to the fault-scaled throughput. Even at a significant 10% failure rate (e.g., 154 failed GPUs for GPT-3 145.6B), ReCycle allows clusters to continue training while observing only between $0.5\%$ and $11.5\%$ degradation from the fault-scaled case. This indicates that ReCycle is well-suited for training tasks in supercomputing-scale clusters with dynamic resource availability (Jiang et al., 2024; He et al., 2023; Zu et al., 2024).
6.4. ReCycle Performance Breakdown
@@
Figure 11. The contribution of the three optimization techniques to ReCycle’s training throughput.
@@
Figure 12. Memory utilization for ReCycle across pipeline stages for GPT-3 6.7B with 30m failure rates compared to DeepSpeed (failure-free). The dashed line shows the 80GB memory capacity of the A100 GPU.
How does each of ReCycle’s techniques contribute to performance?
@@
To assess the benefits and necessity of each technique in ReCycle, as outlined in Section 3, we conducted an ablation study. This involved repeating the real-world experiment on the Azure cluster with GPT-3 models, using a failure frequency of 30 minutes. We adjusted the MILP formulation in the Planner to progressively enable each optimization. Figure 11 shows the throughput achieved at each stage of the study, normalized to the fault-free throughput.
@@
Adaptive Pipelining allows ReCycle to continue training despite failures, but it experiences significant throughput degradation due to the overhead from additional work required by peers of the failed workers. By introducing Decoupled BackProp, ReCycle boosts the training throughput by 63% to 118%, effectively utilizing bubbles during the cool-down phase of the pipeline schedule to mask this overhead (see Figure 3(b)). Further enhancements come from the Staggered Optimizer, which boosts throughput by an additional $7\%$ to $11\%$ by leveraging bubbles from the warm-up phase of the pipeline schedule.
Can ReCycle efficiently exploit unused GPU memory in hybrid parallelism?
@@
We recognize that GPU memory utilization varies across different stages in hybrid parallelism, with later pipeline stages typically requiring less memory. To leverage this, the Decoupled BackProp technique utilizes the available surplus GPU memory as a buffer, allowing for the postponement of $B_{weight}$ computations (see Figure 5). Figure 12 shows the peak GPU memory utilization across the eight stages of pipelined execution for ReCycle compared to the fault-free DeepSpeed, based on training the GPT-3 6.7B model with a 30-minute failure interval.
@@
DeepSpeed – and any system utilizing a pipeline schedule akin to 1F1B – fails to fully utilize GPU memory, especially in the later pipeline stages (Narayanan et al., 2019). This surplus arises because GPU workers in earlier stages must store more intermediate results, as shown in Figure 3. In contrast, ReCycle effectively capitalizes on this opportunity, achieving near-complete utilization of GPU memory to optimize adaptive pipelined execution in the presence of failures. The Planner’s MILP formulation is tailored to model and manage these additional memory requirements while adhering to GPU memory constraints, thereby preventing memory exhaustion.
@@
Figure 13. ReCycle Planner latency (in seconds) to find optimized schedules for up to $25\%$ of failed GPUs.
What is the overhead of ReCycle’s Planner?
@@
To evaluate the latency of the Planner in generating adaptive schedules for large clusters, Figure 13 shows the time required to generate all necessary schedules for up to 25% node failures across various hybrid parallelism strategies. Users have the flexibility to configure the fault-tolerance threshold to suit their needs, whether smaller or larger. The reported latency encompasses both phases of the Planner and was measured using the Gurobi MILP solver (Gurobi, 2024) on a 96-core CPU.
In a training setup with 2048 GPUs (with $DP=32$ and $PP=64$), the Planner can generate all adaptive schedules to accommodate up to $512$ failures in just $3153$ seconds (or $52.5$ minutes). Considering that training large models on thousands of GPUs typically takes weeks (Smith et al., 2022; Jiang et al., 2024; Naumov et al., 2020), the Planner’s latency is negligible, accounting for less than $0.1\%$ of the overall training time.